# Efficient designs for multivariate crossover trials

### v1

@@ Appendix B Efficiency under the generalized markov-type structure
$\begin{split}\bm{\varOmega}_{1}&=\bm{V}_{1}^{*}+\frac{\bar{\rho}^{2}}{\sigma_{12}}\bm{V}_{R}^{*},\\ \bm{\varOmega}_{2}&=\frac{\bar{\rho}}{\sigma_{12}}\bm{V}_{R}^{*},\text{ and}\\ \bm{\varOmega}_{4}&=\frac{1}{\sigma_{12}}\bm{V}_{R}^{*}.\end{split}$ (S.24)
Here, $\bm{H}_{n}=\bm{I}_{n}-\frac{1}{n}\bm{{1}}_{n}\bm{{1}}_{n}^{{}^{\prime}}$, $\bm{V}_{1}^{*}=\bm{V}_{1}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{1}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{1}^{-1}\bm{J}_{p\times p}\bm{V}_{1}^{-1}$, $\bm{V}_{R}^{*}=\bm{V}_{R}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{R}^{-1}\bm{{1}}_{p}\right)^{-1}\times$
$\bm{V}_{R}^{-1}\bm{J}_{p\times p}\bm{V}_{R}^{-1}$, $\bm{J}_{p\times p}=\bm{{1}}_{p}\bm{{1}}_{p}^{{}^{\prime}}$, $\bar{\rho}=\rho\sqrt{\frac{\sigma_{22}}{\sigma_{11}}}$ and $\sigma_{12}=\sigma_{22}\left(1-\rho^{2}\right)$.
@@
$\bm{\varSigma}=\begin{bmatrix}\bm{\varSigma}_{11}&\bm{\varSigma}_{12}\\ \bm{\varSigma}_{21}&\bm{\varSigma}_{22}\end{bmatrix},$ (S.27)
where $\bm{\varSigma}_{11}=\bm{I}_{n}\otimes\bm{V}_{1}$, $\bm{\varSigma}_{12}=\bm{\varSigma}_{21}=\rho\sqrt{\frac{\sigma_{22}}{\sigma_{11}}}\bm{\varSigma}_{11}$, $\bm{\varSigma}_{22}=\rho^{2}\frac{\sigma_{22}}{\sigma_{11}}\bm{\varSigma}_{11}+\sigma_{22}(1-\rho^{2})\bm{\varSigma}_{R}$ and $\bm{\varSigma}_{R}=\bm{I}_{n}\otimes\bm{V}_{R}$. Here, $\bm{V}_{1}$ is a known positive definite and symmetric matrix with each diagonal element as $\sigma_{11}>0$, and $\bm{V}_{R}$ is a known correlation matrix. Also, $\sigma_{11},\sigma_{22}>0$ and $10<|\rho|<$ 1. So,
$\bm{\varSigma}^{-1}$ $=\begin{bmatrix}\bm{\varSigma}^{11}&\bm{\varSigma}^{12}\\ \bm{\varSigma}^{21}&\bm{\varSigma}^{22}\end{bmatrix},$ (S.30)
where $\bm{\varSigma}^{11}=\bm{\varSigma}_{11}^{-1}+\bm{\varSigma}_{11}^{-1}\bm{\varSigma}_{12}\bm{\varSigma}^{22}\bm{\varSigma}_{21}\bm{\varSigma}_{11}^{-1}$, $\bm{\varSigma}^{12}=\left(\bm{\varSigma}^{21}\right)^{{}^{\prime}}=-\bm{\varSigma}_{11}^{-1}\bm{\varSigma}_{12}\bm{\varSigma}^{22}$ and $\bm{\varSigma}^{22}=$
$\left(\bm{\varSigma}_{11}-\bm{\varSigma}_{12}\bm{\varSigma}^{-1}_{22}\bm{\varSigma}_{21}\right)^{-1}$.
@@
$\begin{split}\bm{\varSigma}^{11}&=\bm{I}_{n}\otimes\bm{K}_{1}^{-1},\text{ and}\\ \bm{\varSigma}^{12}&=\bm{\varSigma}^{21}=-\bar{\rho}\bm{\varSigma}^{22}=-\bm{I}_{n}\otimes\bar{\rho}\bm{K}_{2}^{-1}.\end{split}$ (S.31)
where $\bm{K}_{1}^{-1}=\bm{V}_{1}^{-1}+\bar{\rho}^{2}\bm{K}_{2}^{-1}$ and $\bm{K}_{2}=\sigma_{12}\bm{V}_{R}$. Note that $\bm{K}_{1}$ and $\bm{K}_{2}$ are symmetric matrices. Since, $\bm{V}_{R}$ is positive definite matrix, $\sigma_{12}>0$ and $|\bar{\rho}|>0$, we get that $\bm{K}_{2}$ is a positive definite matrix. Since $\bm{\varSigma}$ is positive definite, $\bm{\varSigma}^{11}$ is also a positive definite matrix. Thus $\bm{K}_{1}$ is a positive definite matrix.
Now, we will find $\left(\bm{Z}_{1}^{{}^{\prime}}\bm{\varSigma}^{-1}\bm{Z}_{1}\right)^{-}$. Since from (S.31), $\bm{\varSigma}^{12}=\bm{\varSigma}^{21}$, using (2.6) and (S.30), we get
$\bm{Z}^{{}^{\prime}}_{1}\bm{\varSigma}^{-1}\bm{Z}_{1}$ $=\begin{bmatrix}\bm{X}^{{}^{\prime}}_{1}\bm{\varSigma}^{11}\bm{X}_{1}&\bm{X}^{{}^{\prime}}_{1}\bm{\varSigma}^{12}\bm{X}_{1}\\ \bm{X}^{{}^{\prime}}_{1}\bm{\varSigma}^{12}\bm{X}_{1}&\bm{X}^{{}^{\prime}}_{1}\bm{\varSigma}^{22}\bm{X}_{1}\end{bmatrix}.$
Let $\bm{A}^{(1)}=\bm{X}^{{}^{\prime}}_{1}\bm{\varSigma}^{11}\bm{X}_{1}$, $\bm{B}^{(1)}=\bm{X}^{{}^{\prime}}_{1}\bm{\varSigma}^{12}\bm{X}_{1}$ and $\bm{D}^{(1)}=\bm{X}^{{}^{\prime}}_{1}\bm{\varSigma}^{22}\bm{X}_{1}$. So,
$\left(\bm{Z}^{{}^{\prime}}_{1}\bm{\varSigma}^{-1}\bm{Z}_{1}\right)^{-}$ $=\begin{bmatrix}\bm{L}_{1}&\bm{L}_{2}\\ \bm{L}_{3}&\bm{L}_{4}\end{bmatrix},$ (S.34)
@@
where $\bm{A}^{(2)}=n\bm{K}_{1}^{-1}$, $\bm{B}^{(2)}=\left(\bm{C}^{(2)}\right)^{{}^{\prime}}=\bm{{1}}_{n}^{{}^{\prime}}\otimes\bm{K}_{1}^{-1}\bm{{1}}_{p}$ and $\bm{D}^{(2)}=\delta_{1}^{-1}\bm{I}_{n}$, where $\delta_{1}=1/\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{1}^{-1}\bm{{1}}_{p}$. Note that since $\bm{K}_{1}$ is a positive definite matrix, $\bm{K}_{1}^{-1}$ is a positive definite matrix, and thus $\delta_{1}>0$.
Hence,
@@
Here $\bm{J}_{p\times p}=\bm{{1}}_{p}\bm{{1}}_{p}^{{}^{\prime}}$.
So, using (S.41), (S.40) becomes
@@
where $\delta_{2}=1/\bm{{1}}^{{}^{\prime}}_{p}\bm{K}_{2}^{-1}\bm{{1}}_{p}$. Note that since $\bm{K}_{2}$ is a positive definite matrix, $\bm{K}_{2}^{-1}$ is a positive definite matrix, and thus $\delta_{2}>0$.
Let $\bm{K}_{3}=\bar{\rho}^{2}\bm{V}_{1}+\bm{K}_{2}$. Note that $\bm{K}_{3}$ is a symmetric matrix. Since $\bm{V}_{1}$, $\bm{K}_{1}$ and $\bm{K}_{2}$ are positive definite matrices, using the result that $\bm{I}_{p}+\bm{\zeta}_{3}\bm{\zeta}_{4}$ is invertible if and only if $\bm{I}_{p}+\bm{\zeta}_{4}\bm{\zeta}_{3}$ is invertible, where $\bm{\zeta}_{3}$ and $\bm{\zeta}_{4}$ are matrices such that $\bm{\zeta}_{3}\bm{\zeta}_{4}$ is a $p\times p$ matrix, we get that $\bm{K}_{3}$ is a invertible matrix. Using the result $\left(\bm{I}_{p}+\bm{\zeta}_{3}\bm{\zeta}_{4}\right)^{-1}=\bm{I}_{p}-\bm{\zeta}_{3}\left(\bm{I}_{p}+\bm{\zeta}_{4}\bm{\zeta}_{3}\right)^{-1}\bm{\zeta}_{4}$, where $\bm{\zeta}_{3}$ and $\bm{\zeta}_{4}$ are matrices such that $\bm{\zeta}_{3}\bm{\zeta}_{4}$ is a $p\times p$ matrix, we get that $\bm{K}_{2}^{-1}-\bar{\rho}^{2}\bm{K}_{2}^{-1}\bm{K}_{1}\bm{K}_{2}^{-1}=\bm{K}_{3}^{-1}$. Thus from (S.44) and (S.47), by calculation we get
@@
where $\delta_{4}=\delta_{2}^{-1}-\bar{\rho}^{2}\delta_{1}\delta_{2}^{-2}$. Here, $\bm{J}_{n\times n}=\bm{{1}}_{n}\bm{{1}}_{n}^{{}^{\prime}}$ and $\bm{H}_{n}=\bm{I}_{n}-\frac{1}{n}\bm{{1}}_{n}\bm{{1}}_{n}^{{}^{\prime}}$. Using the expression of $\bm{K}_{1}^{-1}$, $\delta_{1}$ and $\delta_{2}$, we get
@@
Let $\bm{A}^{(3)}=n\bm{K}_{3}^{-1}$, $\bm{B}^{(3)}=\left(\bm{C}^{(3)}\right)^{{}^{\prime}}=\bm{{1}}_{n}^{{}^{\prime}}\otimes\bm{K}_{3}^{-1}\bm{{1}}_{p}$ and $\bm{D}^{(3)}=\delta_{4}\bm{H}_{n}+\frac{1}{n}\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{3}^{-1}\bm{{1}}_{p}\right)\bm{J}_{n\times n}$. So,
@@
$\left(\bm{Z}^{{}^{\prime}}_{1}\bm{\varSigma}^{-1}\bm{Z}_{1}\right)^{-}$ $=\begin{bmatrix}\bm{L}_{1}&\bm{L}_{2}\\ \bm{L}_{3}&\bm{L}_{4}\end{bmatrix},$ (S.67)
@@
Now, we find $\bm{\varSigma}^{-1}\bm{Z}_{1}$. From (2.6), (S.30) and (S.31), by calculation we get
$\bm{\varSigma}^{-1}\bm{Z}_{1}$ $=\begin{bmatrix}\bm{\varSigma}^{11}\bm{X}_{1}&\bm{\varSigma}^{12}\bm{X}_{1}\\ \bm{\varSigma}^{12}\bm{X}_{1}&\bm{\varSigma}^{22}\bm{X}_{1}\end{bmatrix},$ (S.71)
@@
$\begin{split}\bm{\varSigma}^{11}\bm{X}_{1}&=\begin{bmatrix}\bm{{1}}_{n}\otimes\bm{K}_{1}^{-1}&\bm{I}_{n}\otimes\bm{K}_{1}^{-1}\bm{{1}}_{p}\end{bmatrix},\text{ and}\\ \bm{\varSigma}^{12}\bm{X}_{1}&=-\bar{\rho}\bm{\varSigma}^{22}\bm{X}_{1}=\begin{bmatrix}-\bm{{1}}_{n}\otimes\bar{\rho}\bm{K}_{2}^{-1}&-\bm{I}_{n}\otimes\bar{\rho}\bm{K}_{2}^{-1}\bm{{1}}_{p}\end{bmatrix}.\end{split}$ (S.72)
Now, we find $\bm{\varSigma}^{-1}\bm{Z}_{1}\left(\bm{Z}^{{}^{\prime}}_{1}\bm{\varSigma}^{-1}\bm{Z}_{1}\right)^{-}\bm{Z}^{{}^{\prime}}_{1}\bm{\varSigma}^{-1}$. Since $\bm{V}_{1}$, $\bm{K}_{1}$, $\bm{K}_{2}$ and $\bm{K}_{3}$ are symmetric matrices, from (S.60), we get that $\bm{K}_{2}^{-1}\bm{K}_{3}=\bm{K}_{1}^{-1}\bm{V}_{1}$ and $\bm{K}_{1}^{-1}\bm{V}_{1}\bm{K}_{2}^{-1}\bm{K}_{1}=\bm{K}_{2}^{-1}\bm{V}_{1}$. So, using the expression of matrix $\bm{K}_{3}$, we get that $\bm{K}_{2}^{-1}\bm{K}_{3}=\bm{I}_{p}+\bar{\rho}^{2}\bm{K}_{2}^{-1}\bm{V}_{1}$ and $\bm{K}_{2}^{-1}\bm{K}_{1}+\bar{\rho}^{2}\bm{K}_{2}^{-1}\bm{V}_{1}\bm{K}_{2}^{-1}\bm{K}_{1}=\bm{K}_{2}^{-1}\bm{V}_{1}$. Using the expression of $\delta_{4}$, by solving we get $\delta_{1}+\bar{\rho}^{2}\delta_{4}^{-1}\delta_{1}^{2}\delta_{2}^{-2}=\delta_{4}^{-1}\delta_{1}\delta_{2}^{-1}$. Thus from (S.67), (S.68), (S.71) and (S.72), by calculation we get
$\bm{\varSigma}^{-1}\bm{Z}_{1}\left(\bm{Z}^{{}^{\prime}}_{1}\bm{\varSigma}^{-1}\bm{Z}_{1}\right)^{-}\bm{Z}^{{}^{\prime}}_{1}\bm{\varSigma}^{-1}$ $=\begin{bmatrix}\bm{A}^{(4)}&\bm{B}^{(4)}\\ \bm{B}^{(4)}&\bm{D}^{(4)}\end{bmatrix},$ (S.75)
@@
$\bm{B}^{(4)}$ $=-\bar{\rho}\bm{D}^{(4)}=-\frac{1}{n}\bm{J}_{n\times n}\otimes\bar{\rho}\bm{K}_{2}^{-1}+\bm{H}_{n}\otimes\bar{\rho}^{3}\delta_{4}^{-1}\delta_{1}\delta_{2}^{-1}\bm{K}_{2}^{-1}\bm{J}_{p\times p}\bm{K}_{2}^{-1}-\bm{H}_{n}\otimes\bar{\rho}\delta_{4}^{-1}\bm{K}_{2}^{-1}\bm{J}_{p\times p}\bm{K}_{2}^{-1}.$
@@
The matrices $\bm{\varOmega}_{1}$, $\bm{\varOmega}_{2}$, and $\bm{\varOmega}_{4}$ have row sums and column sums as zero.
@@
Using the expression of the matrices $\bm{\varOmega}_{1}$, $\bm{\varOmega}_{2}$, and $\bm{\varOmega}_{4}$, as given in Theorem S.4, we get that $\bm{{1}}_{p}^{{}^{\prime}}\bm{\varOmega}_{1}=\bm{{1}}_{p}^{{}^{\prime}}\bm{\varOmega}_{2}=\bm{{1}}_{p}^{{}^{\prime}}\bm{\varOmega}_{4}=\bm{{0}}_{1\times p}$ and $\bm{\varOmega}_{1}\bm{{1}}_{p}=\bm{\varOmega}_{2}\bm{{1}}_{p}=\bm{\varOmega}_{4}\bm{{1}}_{p}=\bm{{0}}_{p\times 1}$. ∎

### v2

@@ Appendix B Efficiency under the generalized Markov-type structure
$\begin{split}\bm{\mathit{\Omega}}_{1}&=\bm{V}_{1}^{*}+\frac{\bar{\rho}^{2}}{\sigma_{12}}\bm{V}_{R}^{*},\\ \bm{\mathit{\Omega}}_{2}&=\frac{\bar{\rho}}{\sigma_{12}}\bm{V}_{R}^{*},\text{ and}\\ \bm{\mathit{\Omega}}_{4}&=\frac{1}{\sigma_{12}}\bm{V}_{R}^{*}.\end{split}$ (S.16)
Here, $\bm{V}_{1}^{*}=\bm{V}_{1}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{1}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{1}^{-1}\bm{J}_{p\times p}\bm{V}_{1}^{-1}$, $\bm{V}_{R}^{*}=\bm{V}_{R}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{R}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{R}^{-1}\bm{J}_{p\times p}\bm{V}_{R}^{-1}$, $\bar{\rho}=\rho\sqrt{\frac{\sigma_{22}}{\sigma_{11}}}$ and $\sigma_{12}=\sigma_{22}\left(1-\rho^{2}\right)$.
@@
$\bm{\mathit{\Sigma}}=\begin{bmatrix}\bm{\mathit{\Sigma}}_{11}&\bm{\mathit{\Sigma}}_{12}\\ \bm{\mathit{\Sigma}}_{21}&\bm{\mathit{\Sigma}}_{22}\end{bmatrix},$ (S.17)
where $\bm{\mathit{\Sigma}}_{11}=\bm{I}_{n}\otimes\bm{V}_{1}$, $\bm{\mathit{\Sigma}}_{12}=\bm{\mathit{\Sigma}}_{21}=\rho\sqrt{\frac{\sigma_{22}}{\sigma_{11}}}\bm{\mathit{\Sigma}}_{11}$, $\bm{\mathit{\Sigma}}_{22}=\rho^{2}\frac{\sigma_{22}}{\sigma_{11}}\bm{\mathit{\Sigma}}_{11}+\sigma_{22}(1-\rho^{2})\bm{\mathit{\Sigma}}_{R}$ and $\bm{\mathit{\Sigma}}_{R}=\bm{I}_{n}\otimes\bm{V}_{R}$. Here, $\bm{V}_{1}$ is a known positive definite and symmetric matrix with each diagonal element as $\sigma_{11}>0$, and $\bm{V}_{R}$ is a known correlation matrix. Also, $\sigma_{11},\sigma_{22}>0$ and $10<|\rho|<$ 1. So
$\bm{\mathit{\Sigma}}^{-1}$ $=\begin{bmatrix}\bm{\mathit{\Sigma}}^{11}&\bm{\mathit{\Sigma}}^{12}\\ \bm{\mathit{\Sigma}}^{21}&\bm{\mathit{\Sigma}}^{22}\end{bmatrix},$ (S.18)
where $\bm{\mathit{\Sigma}}^{11}=\bm{\mathit{\Sigma}}_{11}^{-1}+\bm{\mathit{\Sigma}}_{11}^{-1}\bm{\mathit{\Sigma}}_{12}\bm{\mathit{\Sigma}}^{22}\bm{\mathit{\Sigma}}_{21}\bm{\mathit{\Sigma}}_{11}^{-1}$, $\bm{\mathit{\Sigma}}^{12}=\left(\bm{\mathit{\Sigma}}^{21}\right)^{{}^{\prime}}=-\bm{\mathit{\Sigma}}_{11}^{-1}\bm{\mathit{\Sigma}}_{12}\bm{\mathit{\Sigma}}^{22}$ and $\bm{\mathit{\Sigma}}^{22}=$
$\left(\bm{\mathit{\Sigma}}_{11}-\bm{\mathit{\Sigma}}_{12}\bm{\mathit{\Sigma}}^{-1}_{22}\bm{\mathit{\Sigma}}_{21}\right)^{-1}$.
@@
$\begin{split}\bm{\mathit{\Sigma}}^{11}&=\bm{I}_{n}\otimes\bm{K}_{1}^{-1},\text{ and}\\ \bm{\mathit{\Sigma}}^{12}&=\bm{\mathit{\Sigma}}^{21}=-\bar{\rho}\bm{\mathit{\Sigma}}^{22}=-\bm{I}_{n}\otimes\bar{\rho}\bm{K}_{2}^{-1}.\end{split}$ (S.19)
where $\bm{K}_{1}^{-1}=\bm{V}_{1}^{-1}+\bar{\rho}^{2}\bm{K}_{2}^{-1}$ and $\bm{K}_{2}=\sigma_{12}\bm{V}_{R}$. Note that $\bm{K}_{1}$ and $\bm{K}_{2}$ are symmetric matrices. Since, $\bm{V}_{R}$ is positive definite matrix, $\sigma_{12}>0$ and $|\bar{\rho}|>0$, we get that $\bm{K}_{2}$ is a positive definite matrix. Since $\bm{\mathit{\Sigma}}$ is positive definite, $\bm{\mathit{\Sigma}}^{11}$ is also a positive definite matrix. Thus, $\bm{K}_{1}$ is a positive definite matrix.
Now, we will find $\left(\bm{Z}_{1}^{{}^{\prime}}\bm{\mathit{\Sigma}}^{-1}\bm{Z}_{1}\right)^{-}$. Since from (S.19), $\bm{\mathit{\Sigma}}^{12}=\bm{\mathit{\Sigma}}^{21}$, using (2.3) and (S.18), we get
$\bm{Z}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{-1}\bm{Z}_{1}$ $=\begin{bmatrix}\bm{X}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{11}\bm{X}_{1}&\bm{X}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{12}\bm{X}_{1}\\ \bm{X}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{12}\bm{X}_{1}&\bm{X}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{22}\bm{X}_{1}\end{bmatrix}.$
Let $\bm{A}^{(1)}=\bm{X}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{11}\bm{X}_{1}$, $\bm{B}^{(1)}=\bm{X}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{12}\bm{X}_{1}$ and $\bm{D}^{(1)}=\bm{X}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{22}\bm{X}_{1}$. So
$\left(\bm{Z}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{-1}\bm{Z}_{1}\right)^{-}$ $=\begin{bmatrix}\bm{L}_{1}&\bm{L}_{2}\\ \bm{L}_{3}&\bm{L}_{4}\end{bmatrix},$ (S.20)
@@
where $\bm{A}^{(2)}=n\bm{K}_{1}^{-1}$, $\bm{B}^{(2)}=\left(\bm{C}^{(2)}\right)^{{}^{\prime}}=\bm{{1}}_{n}^{{}^{\prime}}\otimes\bm{K}_{1}^{-1}\bm{{1}}_{p}$ and $\bm{D}^{(2)}=\delta_{1}^{-1}\bm{I}_{n}$, where $\delta_{1}=1/\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{1}^{-1}\bm{{1}}_{p}$. Note that since $\bm{K}_{1}$ is a positive definite matrix, $\bm{K}_{1}^{-1}$ is a positive definite matrix and thus $\delta_{1}>0$.
Hence
@@
So using (S.23), (S.22) becomes
@@
where $\delta_{2}=1/\bm{{1}}^{{}^{\prime}}_{p}\bm{K}_{2}^{-1}\bm{{1}}_{p}$. Note that since $\bm{K}_{2}$ is a positive definite matrix, $\bm{K}_{2}^{-1}$ is a positive definite matrix and thus $\delta_{2}>0$.
Let $\bm{K}_{3}=\bar{\rho}^{2}\bm{V}_{1}+\bm{K}_{2}$. Note that $\bm{K}_{3}$ is a symmetric matrix. Since $\bm{V}_{1}$, $\bm{K}_{1}$ and $\bm{K}_{2}$ are positive definite matrices, using the result that $\bm{I}_{p}+\bm{\mathit{\zeta}_{3}}\bm{\mathit{\zeta}_{4}}$ is invertible if and only if $\bm{I}_{p}+\bm{\mathit{\zeta}_{4}}\bm{\mathit{\zeta}_{3}}$ is invertible, where $\bm{\mathit{\zeta}_{3}}$ and $\bm{\mathit{\zeta}_{4}}$ are matrices such that $\bm{\mathit{\zeta}_{3}}\bm{\mathit{\zeta}_{4}}$ is a $p\times p$ matrix, we get that $\bm{K}_{3}$ is a invertible matrix. Using the result $\left(\bm{I}_{p}+\bm{\mathit{\zeta}_{3}}\bm{\mathit{\zeta}_{4}}\right)^{-1}=\bm{I}_{p}-\bm{\mathit{\zeta}_{3}}\left(\bm{I}_{p}+\bm{\mathit{\zeta}_{4}}\bm{\mathit{\zeta}_{3}}\right)^{-1}\bm{\mathit{\zeta}_{4}}$, where $\bm{\mathit{\zeta}_{3}}$ and $\bm{\mathit{\zeta}_{4}}$ are matrices such that $\bm{\mathit{\zeta}_{3}}\bm{\mathit{\zeta}_{4}}$ is a $p\times p$ matrix, we get that $\bm{K}_{2}^{-1}-\bar{\rho}^{2}\bm{K}_{2}^{-1}\bm{K}_{1}\bm{K}_{2}^{-1}=\bm{K}_{3}^{-1}$. Thus from (S.24) and (S.25), by calculation we get
@@
where $\delta_{4}=\delta_{2}^{-1}-\bar{\rho}^{2}\delta_{1}\delta_{2}^{-2}$. Using the expression of $\bm{K}_{1}^{-1}$, $\delta_{1}$ and $\delta_{2}$, we get
@@
Let $\bm{A}^{(3)}=n\bm{K}_{3}^{-1}$, $\bm{B}^{(3)}=\left(\bm{C}^{(3)}\right)^{{}^{\prime}}=\bm{{1}}_{n}^{{}^{\prime}}\otimes\bm{K}_{3}^{-1}\bm{{1}}_{p}$ and $\bm{D}^{(3)}=\delta_{4}\bm{H}_{n}+\frac{1}{n}\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{3}^{-1}\bm{{1}}_{p}\right)\bm{J}_{n\times n}$. So
@@
$\left(\bm{Z}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{-1}\bm{Z}_{1}\right)^{-}$ $=\begin{bmatrix}\bm{L}_{1}&\bm{L}_{2}\\ \bm{L}_{3}&\bm{L}_{4}\end{bmatrix},$ (S.35)
@@
Now, we find $\bm{\mathit{\Sigma}}^{-1}\bm{Z_{1}}$. From (2.3), (S.18) and (S.19), by calculation we get
$\bm{\mathit{\Sigma}}^{-1}\bm{Z_{1}}$ $=\begin{bmatrix}\bm{\mathit{\Sigma}}^{11}\bm{X_{1}}&\bm{\mathit{\Sigma}}^{12}\bm{X_{1}}\\ \bm{\mathit{\Sigma}}^{12}\bm{X_{1}}&\bm{\mathit{\Sigma}}^{22}\bm{X_{1}}\end{bmatrix},$ (S.37)
@@
$\begin{split}\bm{\mathit{\Sigma}}^{11}\bm{X_{1}}&=\begin{bmatrix}\bm{{1}}_{n}\otimes\bm{K}_{1}^{-1}&\bm{I}_{n}\otimes\bm{K}_{1}^{-1}\bm{{1}}_{p}\end{bmatrix},\text{ and}\\ \bm{\mathit{\Sigma}}^{12}\bm{X_{1}}&=-\bar{\rho}\bm{\mathit{\Sigma}}^{22}\bm{X_{1}}=\begin{bmatrix}-\bm{{1}}_{n}\otimes\bar{\rho}\bm{K}_{2}^{-1}&-\bm{I}_{n}\otimes\bar{\rho}\bm{K}_{2}^{-1}\bm{{1}}_{p}\end{bmatrix}.\end{split}$ (S.38)
Now, we find $\bm{\mathit{\Sigma}}^{-1}\bm{Z_{1}}\left(\bm{Z}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{-1}\bm{Z}_{1}\right)^{-}\bm{Z}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{-1}$. Since $\bm{V}_{1}$, $\bm{K}_{1}$, $\bm{K}_{2}$ and $\bm{K}_{3}$ are symmetric matrices, from (S.32), we get that $\bm{K}_{2}^{-1}\bm{K}_{3}=\bm{K}_{1}^{-1}\bm{V}_{1}$ and $\bm{K}_{1}^{-1}\bm{V}_{1}\bm{K}_{2}^{-1}\bm{K}_{1}=\bm{K}_{2}^{-1}\bm{V}_{1}$. So, using the expression of matrix $\bm{K}_{3}$, we get that $\bm{K}_{2}^{-1}\bm{K}_{3}=\bm{I}_{p}+\bar{\rho}^{2}\bm{K}_{2}^{-1}\bm{V}_{1}$ and $\bm{K}_{2}^{-1}\bm{K}_{1}+\bar{\rho}^{2}\bm{K}_{2}^{-1}\bm{V}_{1}\bm{K}_{2}^{-1}\bm{K}_{1}=\bm{K}_{2}^{-1}\bm{V}_{1}$. Using the expression of $\delta_{4}$, by solving we get $\delta_{1}+\bar{\rho}^{2}\delta_{4}^{-1}\delta_{1}^{2}\delta_{2}^{-2}=\delta_{4}^{-1}\delta_{1}\delta_{2}^{-1}$. Thus from (S.35), (S.36), (S.37) and (S.38), by calculation we get
$\bm{\mathit{\Sigma}}^{-1}\bm{Z_{1}}\left(\bm{Z}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{-1}\bm{Z}_{1}\right)^{-}\bm{Z}^{{}^{\prime}}_{1}\bm{\mathit{\Sigma}}^{-1}$ $=\begin{bmatrix}\bm{A}^{(4)}&\bm{B}^{(4)}\\ \bm{B}^{(4)}&\bm{D}^{(4)}\end{bmatrix},$ (S.39)
@@
$\bm{B}^{(4)}$ $=-\bar{\rho}\bm{D}^{(4)}$
$=-\frac{1}{n}\bm{J}_{n\times n}\otimes\bar{\rho}\bm{K}_{2}^{-1}+\bm{H}_{n}\otimes\bar{\rho}^{3}\delta_{4}^{-1}\delta_{1}\delta_{2}^{-1}\bm{K}_{2}^{-1}\bm{J}_{p\times p}\bm{K}_{2}^{-1}-\bm{H}_{n}\otimes\bar{\rho}\delta_{4}^{-1}\bm{K}_{2}^{-1}\bm{J}_{p\times p}\bm{K}_{2}^{-1}.$
@@
The matrices $\bm{\mathit{\Omega}}_{1}$, $\bm{\mathit{\Omega}}_{2}$, and $\bm{\mathit{\Omega}}_{4}$ have row sums and column sums as zero.
@@
Using the expression of the matrices $\bm{\mathit{\Omega}}_{1}$, $\bm{\mathit{\Omega}}_{2}$, and $\bm{\mathit{\Omega}}_{4}$, as given in Lemma S.4, we get that $\bm{{1}}_{p}^{{}^{\prime}}\bm{\mathit{\Omega}}_{1}=\bm{{1}}_{p}^{{}^{\prime}}\bm{\mathit{\Omega}}_{2}=\bm{{1}}_{p}^{{}^{\prime}}\bm{\mathit{\Omega}}_{4}=\bm{{0}}_{1\times p}$ and $\bm{\mathit{\Omega}}_{1}\bm{{1}}_{p}=\bm{\mathit{\Omega}}_{2}\bm{{1}}_{p}=\bm{\mathit{\Omega}}_{4}\bm{{1}}_{p}=\bm{{0}}_{p\times 1}$. ∎